OpenStudy (anonymous):

A car completes a journey at an average speed of 40km/h. At what speed must it travel on the return journey if the average speed was 60km/h??

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

the answer must be 120 im not getting it?

OpenStudy (cwrw238):

the answer in the book is 120?

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

i might have misread the question is this correct: on first part of the journey the car travels at 40 km/hr the average speed for the whole journey is 60 km / hr we need to find the average speed for the return journey

OpenStudy (cwrw238):

ok if thats the case lets assume the first part and return distance is 1 km ( it does;t matter what you take here) then first part of journey : 40 = 1/t where t is time taken return part v = 1 / t2 where v = average speed of return and t2 = time taken for whole journey we have 60 = 2 / (t + t2) 40 = 1/ t so t = 1/40 v = 1 / t2 so t2 = 1/v

OpenStudy (cwrw238):